OpenStudy (anonymous):

-18x^2y^-1/2 over 6x^-1/5y....Simplify and express the answer with positive exponents. Evaluate numerical expressions.

OpenStudy (anonymous):

do you need help in learning how to do these?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

when you write the expression as a fraction, the letters with the negative powers ...if they are above the viculum they go below and the sign changes in their pwer

OpenStudy (anonymous):

Ok I understand that part

OpenStudy (anonymous):

if they are below the vinculum you take them aboce and change the sign in their powers

OpenStudy (anonymous):

If you end up with ...eg...y^2 x y^1/2...you add the powers

OpenStudy (anonymous):

Ok, add them...Gotcha

OpenStudy (anonymous):

I can lead you through this question if you wish

OpenStudy (anonymous):

Yes if you could please

OpenStudy (anonymous):

Write the whole thing as numerator over denominator

OpenStudy (anonymous):

\[x^2/x^{-1/5}=x^2x^{1/5}=x^{11/5}\]

OpenStudy (anonymous):

did it

OpenStudy (anonymous):

-18/6= -3/1 = -3

OpenStudy (anonymous):

6 can go into the numbers 18 and 6...cancel

OpenStudy (anonymous):

cancel them out?

OpenStudy (anonymous):

\[y ^{-1/2}/y = y ^{1/2}y =y ^{3/2}\]

OpenStudy (anonymous):

you will get -3 on top and the 6 at the bottom cancels

OpenStudy (anonymous):

alright ok

OpenStudy (anonymous):

Notice x has a negative power underneath

OpenStudy (anonymous):

oh opss.... \[y^{-1/2}/y= 1/y^{1/2}y= 1/y^{3/2}\]

OpenStudy (anonymous):

so in the end it's\[-3x^{11/5}/y^{3/2}\]

OpenStudy (anonymous):

i notice it

OpenStudy (anonymous):

carry x^-1/5 on top...it becomes x^1/5 and it is multiplied by x^2

OpenStudy (anonymous):

I am getting the hang of this, I think that I get it now

OpenStudy (anonymous):

The y^-1/2 goes to the bottom

OpenStudy (anonymous):

u'll manage now ?